15.The following information pertains to Diane Company. Assume that all balance sheet amounts represent both average and ending balance figures and that all sales were on credit. Use this information to answer the question that follow.

Assets
Cash and short-term investments $ 30,000
Accounts receivable (net) 20,000
Inventory 15,000
Property, plant, and equipment   185,000
Total assets $250,000


Liabilities and Stockholders' Equity
Current liabilities $ 45,000
Long-term liabilities 70,000
Stockholders' equity—Common   135,000
Total liabilities and stockholders' equity $250,000


Income Statement
Sales $85,000
Cost of goods sold   45,000
Gross margin $40,000
Operating expenses (15,000)
Interest expenses   (5,000)
Net income $20,000
Number of shares of common stock outstanding 6,000
Market price of common stock $20
Total dividends paid $9,000
Cash provided by operations $30,000


Using the data provided for Diane Company, what is the dividend yield?
Answer #1
Dividend per share = 9000/6000= $1.5
Dividend yield = Dividend per share/Market price of common stock = 1.5/20= 7.5%